OpenStudy (anonymous):

find the derivative of y=sin sqrt(1+x^2)

rishavraj (rishavraj):

do it using chain rule....

OpenStudy (anonymous):

i did but i got stuck at \[y'=(\sin)(1/2)(1+x ^{2})^{-1/2}+(1+x ^{2})^{1/2}(\cos)\]

rishavraj (rishavraj):

hmmm addition sign??? see wht would be derivative of \[\sin(x^2 + 1)\] just asking??????

OpenStudy (anonymous):

you use the product rule

rishavraj (rishavraj):

and hey ur question is \[\sin \sqrt{1 + x^2} \]

rishavraj (rishavraj):

product rule is used in questions like 1) \[x^2 e^x\] 2)\[\sin(x) \tan(x)\]

OpenStudy (anonymous):

yea so i got \[y'=(\sin)(1/2)(1+x^2)^{-1/2}(2x)+(1+x^2)^{1/2}(\cos)\]

OpenStudy (anonymous):

then if im wrong how would you find the derivative of that

rishavraj (rishavraj):

nah nah ...u don't use product rule here.... chain rule..... like u see derivative of sin(x^2 + 1) would be \[2x \cos(x^2 + 1)\]

OpenStudy (anonymous):

well the book has a different answer

rishavraj (rishavraj):

hey its not the answer for the question u asked....

rishavraj (rishavraj):

the question u askd the answer is \[\frac{ x \cos \sqrt{1 + x^2} }{ \sqrt{1 + x^2} }\]

OpenStudy (anonymous):

okay yea that right but how did u get that

rishavraj (rishavraj):

for the question u askd put u = sqr rt 1 + x^2 so y = sin u now \[\frac{ dy }{ dx } = \frac{ dy }{ du } \times \frac{ du }{ dx }\]

OpenStudy (anonymous):

so how would you set that up

rishavraj (rishavraj):

if \[u = \sqrt{1 + x^2}\] then wht would be \[\frac{ du }{ dx }\] ????????

OpenStudy (anonymous):

never mind i got it thank you

rishavraj (rishavraj):

u r most welcome....